OpenStudy (anonymous):

With exponential functions, the formula is f(x) = b^(x-h) + k The h moves it left or right. The k moves it up or down. If h is negative, which way will it go: left or right? If k is negative, which way will it go: up or down?

OpenStudy (anonymous):

Okay, so if the function were f(x) = 11^x And the function is moved right 5 and down 8 the new function g(x) = 11^(x+5) - 8 ????

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

What do you mean?

OpenStudy (primeralph):

Talking to yourself?

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

No. He said something. he must've deleted the comment DX

OpenStudy (anonymous):

He said that 11^x was completely out of the realm. And I don't see how

OpenStudy (anonymous):

Ugh, can someone please just explain DX

OpenStudy (mathmate):

When a function is expressed in the form f(x)=a^(x-h)+k h and k represent translations to the right and to the top respectively. Evidently, if any of the two is negative, the direction is reversed.

OpenStudy (anonymous):

wait. so when it is (x-h) it goes right, and when it is (x+h) it goes left?

OpenStudy (primeralph):

- right + left

OpenStudy (primeralph):

Think of is as moving the graphing paper away while keeping the curve still.

OpenStudy (anonymous):

is it that way with the k too?

OpenStudy (mathmate):

Yes, if h and k are both positive.

OpenStudy (mathmate):

The equivalent statement of what you just interpreted would be: translation to the right if h is positive, and to the left if h is negative.

OpenStudy (anonymous):

Primer just said the opposite of what you just said DX

OpenStudy (primeralph):

@SavannahWillett , Me and @mathmate are saying the same thing.

OpenStudy (primeralph):